OpenStudy (anonymous):

Small buttons cost $7 per pound. Large buttons cost $9 per pound. How many pounds of large buttons are needed for an 8–pound mixture that costs $8.50 per pound?

hero (hero):

Hint: x + y = 8 7x + 9y = 8.50(8)

hero (hero):

x = small buttons y = large buttons